Find the Fourier series for the function
f(x) = (x-x)^2,-L < x < L
1
Expert's answer
2020-08-03T18:24:24-0400


"f(x)=x-x^2"

"a_0={1\over 2L}\displaystyle\int_{-L}^Lf(x)dx={1\over 2L}\displaystyle\int_{-L}^L(x-x^2)dx="

"={1\over 2L}[{x^2 \over 2}-{x^3 \over 3}]\begin{matrix}
 L \\
 -L
\end{matrix}={1\over 2L}({L^2 \over 2}-{L^3 \over 3}-({L^2 \over 2}+{L^3 \over 3}))="

"=-{L^2 \over 3}"

"a_n={1\over L}\displaystyle\int_{-L}^Lf(x)\cos ({n\pi x \over b})dx="

"={1\over L}\displaystyle\int_{-L}^L(x-x^2)\cos ({n\pi x \over L})dx"

"\int x\cos({n\pi x \over L})dx={L \over n\pi}x\sin({n\pi x \over L})-{L \over n\pi}\int \sin({n\pi x \over L})dx="

"={L \over n\pi}x\sin({n\pi x \over L})+{L^2 \over n^2\pi^2}\cos({n\pi x \over L})+C_1"


"\int x^2\cos({n\pi x \over L})dx={L \over n\pi}x^2\sin({n\pi x \over L})-{2L \over n\pi}\int x\sin({n\pi x \over L})dx="

"={L x^2\over n\pi}\sin({n\pi x \over L})+{2L^2x \over n^2\pi^2}\cos({n\pi x \over L})-{2L^2 \over n^2\pi^2}\int \cos({n\pi x \over L})dx="

"={Lx^2 \over n\pi}\sin({n\pi x \over L})+{2L^2x \over n^2\pi^2}\cos({n\pi x \over L})-{2L^3 \over n^3\pi^3}\sin({n\pi x \over L})+C_2"


"a_n={1\over L}\bigg[{L \over n\pi}x\sin({n\pi x \over L})+{L^2 \over n^2\pi^2}\cos({n\pi x \over L})\bigg]\begin{matrix}
 L\\
 -L
\end{matrix}-"

"-{1\over L}\bigg[{Lx^2 \over n\pi}\sin({n\pi x \over L})+{2L^2x \over n^2\pi^2}\cos({n\pi x \over L})-{2L^3 \over n^3\pi^3}\sin({n\pi x \over L})\bigg]\begin{matrix}
 L\\
 -L
\end{matrix}="

"=0-(-1)^n {4L^2 \over n^2\pi^2}=-(-1)^n {4L^2 \over n^2\pi^2}"


"b_n={1\over L}\displaystyle\int_{-L}^Lf(x)\sin ({n\pi x \over b})dx="

"={1\over L}\displaystyle\int_{-L}^L(x-x^2)\sin ({n\pi x \over L})dx"

"\int x\sin({n\pi x \over L})dx=-{L \over n\pi}x\cos({n\pi x \over L})+{L \over n\pi}\int \cos({n\pi x \over L})dx="

"=-{L \over n\pi}x\cos({n\pi x \over L})+{L^2 \over n^2\pi^2}\sin({n\pi x \over L})+C_3"

"\int x^2\sin({n\pi x \over L})dx=-{L \over n\pi}x^2\cos({n\pi x \over L})+{2L \over n\pi}\int x\cos({n\pi x \over L})dx="

"=-{L x^2\over n\pi}\cos({n\pi x \over L})+{2L^2x \over n^2\pi^2}\sin({n\pi x \over L})-{2L^2 \over n^2\pi^2}\int \sin({n\pi x \over L})dx="

"=-{L x^2\over n\pi}\cos({n\pi x \over L})+{2L^2x \over n^2\pi^2}\sin({n\pi x \over L})+{2L^3\over n^3\pi^3}\cos({n\pi x \over L})+C_4"


"b_n={1\over L}\bigg[-{Lx \over n\pi}\cos({n\pi x \over L})+{L^2 \over n^2\pi^2}\sin({n\pi x \over L})\bigg]\begin{matrix}
 L\\
 -L
\end{matrix}-"

"-{1\over L}\bigg[-{Lx^2 \over n\pi}\cos({n\pi x \over L})+{2L^2x \over n^2\pi^2}\sin({n\pi x \over L})+{2L^3 \over n^3\pi^3}\cos({n\pi x \over L})\bigg]\begin{matrix}
 L\\
 -L
\end{matrix}="


"=-{2L \over n\pi}(-1)^n-0=-(-1)^n{2L \over n\pi}"


"f(x)=-{L^2 \over 3}+\displaystyle\sum_{i=1}^n(-(-1)^n {4L^2\over n^2\pi^2})\cos({n\pi x \over L}) +"

"+\displaystyle\sum_{i=1}^n(-(-1)^n{2L \over n\pi})\sin({n\pi x \over L})"